Use the distance formula to show that each set of points describes an isosceles triangle
(a triangle with two sides of equal length).
(-2, 4), (2, 7), and (5, 3)

Answer by checkley71(8403)

d=sqrt[(7-4)^2+(2+2)^2]
d=sqrt3^2+4^2
d=sqrt9+16
d=sqrt25
d=5 length of one side
d=sqrt[(3-7)^2+(5-2)^2]
d=sqrt-4^2+3^2
d=sqrt16=9
d=sqrt25
d=5 length of a second side
d=sqrt[(5+2)^2+(3-4)^2
d=sqrt7^2+(-1)^2
d=sqrt49+1
d=sqrt50
d=7.07 third side not equal.
answer: this is an isosceles triangle.